OpenStudy (anonymous):

Can anyone please guide me step by step on: "use the maclaurin series for f(x)=sin^2 Hint: using the identity sin^2 x = 1/2 (1-cos(2x))"

OpenStudy (anonymous):

Follow the hint that \(\large \sin^2x=\frac{1}{2}(1−\cos(2x))\) along with the fact that \[\large \cos u=\sum_{n=0}^{\infty}\frac{(−1)^nu^{2n}}{(2n)!}\tag {1}\]and hence you get the power series for \(\large \cos(2x)\) by letting \(\large u=2x\) in \(\large(1)\). Can you take things from here? :-)

OpenStudy (anonymous):

Thanks. cos U you just put in be course the derivative of sin is cos? so you put that outside?…how does the 1 turn into (-1)?

OpenStudy (anonymous):

There's no differentiating being done here. What I meant by what I wrote originally was that you can rewrite \[\large \sin^2 x = \frac{1}{2}(1-\cos(2x)) = \frac{1}{2}\left(1-\sum_{n=0}^{\infty}\frac{(-1)^n(2x)^{2n}}{(2n)!} \right) = \ldots\tag{1}\]using the trig identity they provided and the fact that we know the Maclaurin series for \(\large \cos x\). All you need to do now is to write \(\large (1)\) as a single power series. Does this clarify things? :-)

OpenStudy (anonymous):

yes i guess so. I have to try out some more of these types. My brain is not that sharp. but i think i got it thanks